OpenStudy (spring98):

The graph shows the location of Jacob's, Kent's, and Hakim's houses. Each unit on the graph represents 1 kilometer.

OpenStudy (spring98):

Hakim is preparing for a cross country meet. Hakim is going to run from his house to Jacob's, then to Kent's, and finally back home. How many total kilometers did Hakim run? Round your answer to the nearest whole kilometer. 6 kilometers 14 kilometers 18 kilometers 22 kilometers

OpenStudy (spring98):

@surjithayer plz help

OpenStudy (anonymous):

it asks for the perimeter of the triangle.

OpenStudy (spring98):

yes and i counted out 12 but there is no 12 on the answer choices

OpenStudy (anonymous):

since it's a right triangle you know that to get the hypotenuse its going to be \[c = \sqrt{a^2+b^2}\] then you'll just add the straight line distances

OpenStudy (anonymous):



OpenStudy (anonymous):

or you could say \[P = A + B + \sqrt{A^2+B^2}\] where A and B are the straight lines

OpenStudy (spring98):

@surjithayer plz help

OpenStudy (anonymous):

\[P = 4 + 4 + \sqrt{4^2 + 4^2}\]

OpenStudy (anonymous):

perimeter=4+4+4*1.414=8+5.656=13.656=14 approx

OpenStudy (anonymous):

13.656 round it up and it's 14

OpenStudy (spring98):

ok so it would be B?

OpenStudy (spring98):

right?

OpenStudy (anonymous):

yes

OpenStudy (spring98):

thanks!!!

OpenStudy (anonymous):

np

OpenStudy (spring98):

i got 100% thank you guys!!!!!! :) :) :)

OpenStudy (anonymous):

gratz